OpenStudy (twistrose):

Susan is creating a solution. A formula for the solution calls for 7/10 mL of water. Susan wants to make 2/7 less of the solution than the formula creates. How much water should Susan use? Express your answer as a fraction in simplest form.

OpenStudy (faiqraees):

So susan wants to make 5/7 of the solution of the formula right?

OpenStudy (twistrose):

yeah

OpenStudy (calculusxy):

I think that we can make a proportion in here. \[\large \frac{ 1 }{ \frac{7}{10} } = \frac{ 1 - \frac{2}{7} }{ x }\] 1 represents the total solution 7/10 is the amount of water required for that total solution 1 - 2/7 shows that she wants to make 2/7 less than the total solution x is the amount of water that you are looking for for the 1- 2/7 mL of solution

OpenStudy (faiqraees):

Okay so the solution of the formula create 7/10ml SO 5/7 of the solution of the formula = 5/7 x 7/10

OpenStudy (calculusxy):

@Twistrose do you understand the proportion that I made?

OpenStudy (twistrose):

no i dont understand that

OpenStudy (faiqraees):

Do you understand my method?

OpenStudy (twistrose):

yeah i understand that better

OpenStudy (faiqraees):

Good @calculusxy method is just a complicated version of my method.

OpenStudy (twistrose):

oh ok

OpenStudy (faiqraees):

Okay so multiply 5/7 by 7/10 what do you get?

OpenStudy (twistrose):

35/70

OpenStudy (faiqraees):

Now simplify it

OpenStudy (calculusxy):

Yes in a way it is:  which is the same thing as \(\large \frac{7}{10} \times \frac{5}{7}\)